OpenStudy (anonymous):

If x=-1.8 and y=2.5 then -x+2y=

OpenStudy (anonymous):

-1.8 + 2*2.5 = -1.8 + 5.0 = 3.2

OpenStudy (anonymous):

If m = -7 and n = 10 then -2m - 3n =